OpenStudy (anonymous):

what is the LCM of 25 and 70?

OpenStudy (anonymous):

25 = 5*5 70 = 7*5*2 lcm = 2*5^2 = 50

OpenStudy (anonymous):

2*5^2*7

OpenStudy (anonymous):

multiples of 25 = 25, 50, 75, 100, 125, 150, 175,200,225, 250, 275, 300, 325, 350,............ multiples of 70 = 70, 140, 210, 280, 350, ......

OpenStudy (anonymous):

therefore the least common multiple is 350

OpenStudy (anonymous):

2*5^2*7 = 350, sorry